OpenStudy (anonymous):

Prove the statement If f'(x) = g'(x) for all x and f(5)=g(5), then f(x) = g(x) for all x for all x>0 Use one of the following: the mean value theorem, Rolle's theorem, the increasing function theorem, the constant function theorem, or the racetrack principle.

OpenStudy (anonymous):

if \[f'(x)=g'(x)\] then \[f(x)=g(x)+c\] and \[f(5)=g(5)\implies c = 0\]

OpenStudy (anonymous):

oh i see what you need, you need the first line that i wrote, that is \[f'(x)=g'(x)\implies f(x)=g(x)+c\] we need to prove that one

OpenStudy (anonymous):

suppose \[h'(x)=0\] for all x, then by the mean value theorem \[\frac{h(x)-h(y)}{x-y}=0\] for all pairs x and y that means for all pairs x and y \[h(x)=h(y)\] showing that \[h(x)=c\] a constant

OpenStudy (anonymous):

apply this to the function \[h(x)=f(x)-g(x)\] since \[f'(x)=g'(x)\] this means that \[h'(x)=f'(x)-g'(x)=0\] for all x, meaning that \[h(x)=f(x)-g(x)=c\] a constant and so \[f(x)=g(x)+c\] for some constant c

OpenStudy (anonymous):

now c is a constant (does not depend on x) we know that \[f(5)=g(5)+c\] and we also are given that \[f(5)=g(5)\] we know that \[c=0\]